# CleanPose: Category-Level Object Pose Estimation via Causal Learning and Knowledge Distillation

### v1

@@ 4.3 Category Knowledge Distillation
where $\psi\in\mathbb{R}^{C_{1}\times C_{3}}$ is a learnable embedding layer and $B$ denotes the batch size. In this way, the tuned point cloud features of our model can effectively receive supervision from generalized ones, facilitating comprehensive category knowledge learning and enhancing the model’s intra-category generalization.
@@ 4.4 Pose Estimation and Overall Loss Function
where $\alpha_{1},\alpha_{2}$ are hyper-parameters to balanced the contribution of each term. We omit some loss terms for for brevity. Please see supplementary for more details.
@@ 5 Experiments
Methods Venue/Source $IoU_{75}^{*}$$\uparrow$ 5°2$\uparrow$ 5°5$\uparrow$ 10°2$\uparrow$ 10°5$\uparrow$
SPD[43] ICCV’20 27.5 19.3 21.4 43.2 54.1
DualPoseNet[21] ICCV’21 30.8 29.3 35.9 50.0 66.8
GPV-Pose[5] CVPR’22 23.1 32.5 43.3 58.2 76.6
6D-ViT[67] TIP’22 33.3 38.2 41.9 59.1 67.9
@@
HS + GeoReF[63] CVPR’24 54.3 51.7 59.6 74.3 83.3
SPD + GeoReF[63] CVPR’24 51.8 54.4 60.3 71.8 79.4
MH6D[29] TNNLS’24 54.2 53.0 61.1 72.0 82.0
@@ 5.1 Comparison with State-of-the-Art Methods
As shown in Fig. 4, our method outperforms existing sota methods across multiple key metrics in different datasets.
Performance on REAL275 dataset. The comparisons between proposed CleanPose and previous methods on challenging REAL275 dataset are shown in Tab. 1. As can be easily observed, our CleanPose achieves the state-of-the-art performance in all metrics and outperforms all previous methods on REAL275 dataset. Significantly, we achieve the precision of 61.5%, 67.4% and 78.3% in the rigorous metric of 5°2 cm, 5°5 cm and 10°2 cm, surpassing the current state-of-the-art method AG-Pose [^1] Throughout the comparison in all experiments, we use the result suggested by the AG-Pose [25] official implementation, which is higher than reported result in the original paper. [25] with a large margin by 4.5%, 2.8% and 3.2%, respectively. Moreover, the qualitative results of AG-Pose and proposed CleanPose are shown in Fig. 5. It can be observed that our method achieves significantly higher precision. These exceptional outcomes further support the efficacy of our approach.
Figure 4: Comparison of CleanPose with sota methods.
@@
Performance on CAMERA25 dataset. The results comparison of our method and the state-of-the-art is presented in Tab. 2. From the observation of results, it can be deduced that CleanPose ranks top performance across all the metrics except $IoU_{75}$, in which our method also achieves comparable performance with the best (80.7% vs. 81.2%). In detail, the proposed CleanPose achieves 80.3% on 5°2 cm, 84.2% on 5°5 cm, 87.7% on 10°2 cm and 92.7% on 10°5 cm, respectively. This superior performance on synthetic dataset further proves the effectiveness of our method.
@@
Queue FIFO 61.5 67.4 78.3 86.2
Queue w/o update 57.1 63.1 77.5 85.7
Queue Similarity 59.1 65.8 77.6 86.0
Memory bank w/o update 57.6 64.4 75.6 83.8
Memory bank Similarity 58.1 65.4 77.5 86.1
@@
5°2cm 5°5cm 10°2cm 10°5cm
w/o fusion 58.5 65.2 79.0 86.5
w/ fusion 61.5 67.4 78.3 86.2
(b) Effect of adaptive weight fusion in front-door adjustment.
@@
Contrastive 55.3 63.1 76.2 86.0
MLP 56.4 63.1 77.6 85.9
MLP + residual 61.5 67.4 78.3 86.2
(c) Effect of distinct feature distillation policies.
@@
PointNet++[40] 57.6 65.3 75.0 84.4
PointMLP[33] 57.0 64.3 77.4 86.2
PointBert[56] 61.5 67.4 78.3 86.2
@@
Performance on HouseCat6D dataset. In Tab. 3, we evaluate our method on HouseCat6D [17] dataset. The proposed CleanPose again achieves state-of-the-art performance in all metrics. In detail, our method outperforms the current best method AG-Pose [25] by 2.9% on $IoU_{50}$, 1.1% on 5°2 cm, 2.0% on 5°5 cm and 2.1% on 10°5 cm, respectively. The overall and category-wise evaluation of 3D IoU on HouseCat6D dataset is provided in supplementary material.
5.2 Ablation Studies
We conduct ablation experiments to demonstrate the effectiveness of the proposed method on the REAL275 dataset [48].
Effect of different feature storage methods and update strategies. In Tab. 5(a), we investigate the impact of distinct combinations of feature storage and update strategies. “Similarity” represents updating the closest features via similarity computation. The results indicate that employing the dynamic queue and FIFO update mechanism yields the best performance. We owe the advantage to the dynamic queue’s superior ability to capture feature variations compared to a memory bank. Moreover, the FIFO mechanism ensures the removal of outdated features, which is beneficial for causal inference.
Effect of adaptive weight fusion method. As shown in Tab. 5(b), our model achieves sota performance even without feature fusion. Introduction of the adaptive feature fusion method further improves the results in rigorous metrics.
Effect of distinct feature distillation policies. In Tab. 5(c), we assess the impact of distinct feature distillation policies. The contrastive learning minimizes the feature distance via feature alignment. Our proposed residual distillation policy achieves the best performance since the residual layers effectively balance feature learning and knowledge transfer.
Effect of different 3D encoders. Tab. 5(d) illustrates the impact of using different 3D encoders in the distillation. Objectively, the pre-trained PointBert [56] realizes the sota point cloud zero-shot classification results [53]. Our model is also more performant using PointBert. It is noteworthy that using PointNet++ [40], which serves as the feature extractor in our model, dose not achieve the best performance. This further demonstrates that our proposed knowledge distillation network focuses on category knowledge rather than feature similarity. Please see supplementary for more ablations.
@@ 6 Conclusion
In this paper, we present CleanPose, the first solution that addresses the confounders in category-level pose estimation from the perspective of causal learning. Motivated by pivotal observation that humans can learn inherent causality beyond biases, we develop a causal inference framework based on front-door adjustment. We formulate the modeling of crucial causal variables in pose estimation task. Guided by the causal modeling, we devise an effective data sampling method and a adaptive feature fusion module to achieve unbiased estimation. Furthermore, we propose a residual knowledge distillation network to transfer rich 3D category information, providing comprehensive category-specific guidance. Extensive experiments on challenging benchmarks REAL275, CAMERA25 and HouseCat6D demonstrate that CleanPose can significantly improve performance, showing the effectiveness of our method.
@@
A Limitation and Broader Impact
B More Loss Function Details
The backbone of our method is based on AG-Pose [25]. In addition to $\mathcal{L}_{pose}$ (Eq. 21), there are some additional loss functions to balance keypoints selection and pose prediction. First, to encourage the keypoints to focus on different parts, the diversity loss $\mathcal{L}_{div}$ is used to force the detected keypoints to be away from each other, in detail:
@@ B More Loss Function Details
Hence, the complete form of overall loss (Eq. 22) is as follows:
@@
C More Details of Using ULIP-2
ULIP-2 [53] is a large-scale 3D foundation model with strong perception capabilities for the point cloud modality. It offers multiple pre-trained versions of point cloud encoders. In our model, there are two key steps that involve the use of different pre-trained encoders of ULIP-2. (i) In the knowledge distillation process, we leverage the pre-trained PointBert [56], which achieves the best zero-shot classification performance across all versions. Therefore, it can provide comprehensive category knowledge guidance for our model. In the ablation study Tab. 5(d), we also compared it with PointNet++ [40], which is more similar in architecture to our model. The experimental results demonstrate that our distillation method focuses more on category knowledge rather than feature similarity. (ii) However, during the initial construction of the dynamic queue, we use the pre-trained PointNet++ [40], as the front-door adjustment primarily focuses on the differences between samples. We aim to avoid introducing confounders due to feature discrepancies from different encoders. The additional ablation study results in Tab. 7(b) also support our analysis.
@@ C More Details of Using ULIP-2
Table 6: Additional ablation studies on some hyper-parameters. Default settings are colored in gray.
D Additional Ablations
Effect of different sampling quantities $N_{s}$. In Sec. 4.2 of main manuscript, we sample $N_{s}$ features for the specific network design to perform front-door adjustment. We study the influence with response to different sampling quantities $N_{s}$ in Tab. 6(b). We find that a large $N_{s}$ leads to a slight performance degradation as selected features may not capture the representation space properly. The results demonstrate that $N_{s}=12$ yields the most significant performance gains.
@@ D Additional Ablations
PointNet++[40] 61.5 67.4 78.3 86.2
PointMLP[33] 56.1 64.8 76.0 85.2
PointBert[56] 58.8 65.2 77.9 86.4
@@
Random 61.5 67.4 78.3 86.2
K-means 58.5 65.1 78.0 86.5
K-means (simi) 56.7 64.2 76.7 86.0
@@
Table 7: Additional ablation studies on confounders queue. Default settings are colored in gray.
Figure 6: Illustration of different feature selectors in ablations Tab. 7(c).
Category $IoU_{25}^{*}$ $IoU_{50}^{*}$ $IoU_{75}^{*}$ 5°2cm 5°5cm 10°2cm 10°5cm
bottle 51.3 49.4 36.9 75.3 81.7 79.9 87.8
bowl 100.0 100.0 93.8 92.8 97.8 95.0 99.9
camera 90.9 83.5 39.5 2.8 3.1 33.9 40.5
can 71.3 71.1 43.1 84.2 85.8 96.9 98.6
laptop 86.3 84.0 76.1 68.9 90.2 71.9 98.5
mug 99.6 99.4 86.1 45.3 45.6 91.9 91.9
Average 83.3 81.2 62.6 61.5 67.4 78.3 86.2
Table 8: Category-wise evaluation of CleanPose on REAL275 dataset. ‘*’ denotes CATRE [31] IoU metrics.
@@
Effect of different 3D encoder for initial construction of the queue. Tab. 7(b) ablates the different 3D encoders of ULIP-2 [53] for initial construction of the dynamic queue. The results exhibit that using PointNet++ [40] yields the most performance gains. As mentioned in Sec. C, the dynamic queue is utilized in the cross-attention phase of front-door adjustment, thus primarily focusing on the differences between samples. Using encoders with similar architectures helps avoid introducing extra confounders.
Various feature selection strategies. In Sec. 4.2 of main manuscript, we randomly sample $N_{s}$ features from queue by default, as shown in Fig. 6(a). Optionally, we can first use K-means to cluster the features of the queue, and then select features from each cluster to form $\mathcal{F}_{samp}$, as illustrated in Fig. 6(b). For fair comparison, the number of clusters is set equal to $N_{s}$. We investigate the impact of these two feature selector in Tab. 7(c). As shown in the table, we observe that K-means-based feature sampling strategy shows a decline in performance on strict metrics (5°2cm and 5°5cm). We argue that k-means, which clusters by Euclidean distance, may lose important boundary information, thus affecting the model performance. Moreover, k-means clustering needs to be performed after each queue update, which increases computational load and training costs. Therefore, we added a comparative experiment using similarity-based updates, denoted as ‘K-means (simi)’, where (simi) refers to the ‘Similarity’ defined in Tab. 5(a). In this case, clustering is only performed once during the initial training. However, experimental results also show that such strategy leads to further performance degradation as one clustering loses the diversity of features.
E More Experimental Results
@@
We report category-wise results of REAL275 [48] in Tab. 8. Since there is a small mistake in the original evaluation code of NOCS [48] for the 3D IoU metrics, we present more reasonable CATRE [31] metrics following [63, 29, 4]. Further, more detailed results of HouseCat6D [17] are shown in Tab. 9. As for more restricted metric $IoU_{75}$, our method also demonstrates state-of-the-art performance (53.9%), further validating the effectiveness of CleanPose in 3D IoU evaluation. Moreover, in category-wise validation on $IoU_{25}$ and $IoU_{50}$, our approach obtains state-of-the-art (e.g., Can, Cup, Glass and Shoe) or competitive results across all categories. It is worth mentioning that our method exhibits more stable performance on these two metrics. For instance, compared to the current sota method AG-Pose [25] in the Box category, our method achieves the best performance (28.7%) on $IoU_{50}$ metric when both obtain competitive results on $IoU_{25}$ metric, with a significant reduction of the AG-Pose (13.9%).

### v2

@@ 4.3 Knowledge Distillation
where $\psi\in\mathbb{R}^{C_{1}\times C_{3}}$ is a learnable embedding layer and $B$ denotes the batch size. In this way, we effectively transfer the debiasing knowledge from the foundation model to the COPE model, further mitigating the interference of data biases. Moreover, the tuned point cloud features of our model can effectively receive supervision from generalized ones, facilitating comprehensive category knowledge learning and enhancing the model’s intra-category generalization.
@@ 4.4 Pose Estimation and Overall Loss Function
where $\alpha_{1},\alpha_{2}$ are hyper-parameters to balanced the contribution of each term. We omit some loss terms for brevity. Please refer to appendix for more details.
@@ 5 Experiments
Methods Venue/Source Shape Prior $IoU_{75}^{*}$$\uparrow$ 5°2$\uparrow$ 5°5$\uparrow$ 10°2$\uparrow$ 10°5$\uparrow$
DPDN[19] ECCV’22 54.0 46.0 50.7 70.4 78.4
MH6D[26] TNNLS’24 54.2 53.0 61.1 72.0 82.0
[…]
@@
SecondPose[3] CVPR’24 49.7 56.2 63.6 74.7 86.0
@@ 5.1 Comparison with State-of-the-Art Methods
As shown in Fig. 4, our method outperforms prior-free SOTA methods across key metrics in different datasets.
Performance on REAL275 dataset. The comparisons between proposed CleanPose and previous methods on challenging REAL275 dataset are shown in Tab. 1. As can be easily observed, our CleanPose achieves the state-of-the-art performance in all metrics and outperforms all previous methods on REAL275 dataset. Significantly, we achieve the precision of 61.7%, 67.6% and 78.3% in the rigorous metric of 5°2 cm, 5°5 cm and 10°2 cm, surpassing the current prior-free state-of-the-art method AG-Pose [22] with a large margin by 4.7%, 3.0% and 3.2%, respectively. Our method even outperforms prior-based sota method GCE-Pose [17] by 4.7%, 2.5% and 2.7% in above three rigorous metric. Moreover, the qualitative results of AG-Pose and proposed CleanPose are shown in Fig. 5. It can be observed that our method achieves significantly higher precision. These exceptional outcomes further support the efficacy of our approach. Please refer to appendix for detailed comparison with SOTA method.
Figure 4: Comparison of ours with prior-free SOTA methods.
@@
Performance on CAMERA25 dataset. The comparison results are presented in Tab. 2. From the observation of results, it can be deduced that CleanPose ranks top performance across all the metrics except $IoU_{75}$, in which our method also achieves comparable performance with the best (80.7% vs. 81.2%). In detail, CleanPose achieves 80.3% on 5°2 cm, 84.2% on 5°5 cm and 87.7% on 10°2 cm, respectively. This superior performance on synthetic dataset further proves the effectiveness of CleanPose.
@@
Performance on HouseCat6D dataset. In Tab. 3, we evaluate our method on HouseCat6D [15] dataset. The proposed CleanPose again achieves state-of-the-art performance in all metrics. In detail, our method outperforms the current best method AG-Pose [22] by 2.9% on $IoU_{50}$, 1.1% on 5°2 cm, 2.0% on 5°5 cm and 2.1% on 10°5 cm, respectively. The overall and category-wise evaluation of 3D IoU on HouseCat6D dataset is provided in appendix.
5.2 Ablation Studies
We conduct ablation experiments to demonstrate the effectiveness of the proposed method on REAL275 dataset [42].
Effect of different feature storage methods and update strategies. In Tab. 5(a), we investigate the impact of distinct combinations of feature storage and update strategies. ‘Similarity’ represents updating the closest features via similarity computation. The results indicate that employing the dynamic queue and FIFO update mechanism yields the best performance. We owe the advantage to the dynamic queue’s superior ability to capture feature variations compared to a memory bank. Moreover, the FIFO ensures the removal of outdated features, which is beneficial for causal inference.
Effect of adaptive fusion method. As shown in Tab. 5(b), our model achieves sota performance even without feature fusion. Introduction of the adaptive feature fusion method further improves the results in rigorous metrics.
Effect of distinct category knowledge learning policies. In Tab. 5(c), we assess the impact of category knowledge learning policies. ‘Concat’ means directly concatenating $\mathcal{F}^{ULIP}_{P}$ and $\mathcal{F}_{P}$, while ‘Contrastive’ indicates aligning above two features via contrastive learning. Our proposed residual distillation policy achieves the best performance since the residual layers effectively balance feature learning and category knowledge transfer.
@@ 5.2 Ablation Studies
Queue FIFO 61.7 67.6 78.3 86.3
Queue w/o update 57.1 66.1 77.5 85.7
Queue Similarity 61.0 67.1 77.6 86.0
Memory bank w/o update 57.6 66.4 75.6 85.5
Memory bank Similarity 59.1 66.9 77.5 86.1
@@
AF 5°2cm 5°5cm 10°2cm 10°5cm
60.5 66.5 78.8 86.5
61.7 67.6 78.3 86.3
(b) Effect of adaptive fusion in front-door adjustment.
@@
Concat 60.0 66.7 77.9 85.9
Contrastive 59.8 66.5 77.6 85.8
Distillation 61.7 67.6 78.3 86.3
(c) Effect of distinct category knowledge learning policies.
@@
PointNet++[35] 59.6 66.8 77.7 85.8
PointMLP[30] 60.8 67.2 77.4 86.0
PointBert[50] 61.7 67.6 78.3 86.3
@@
Effect of different 3D encoders. Tab. 5(d) illustrates the impact of using different 3D encoders in the distillation. Objectively, the pre-trained PointBert [50] realizes the sota point cloud zero-shot classification results [47]. Our model is also more performant using PointBert. Moreover, since our method selects PointNet++ as 3D encoder, we hypothesize that when the teacher and student models share similar architectures (e.g., both using PointNet++), the distillation may mislead the student model to focus on feature structure similarity rather than transferring category knowledge. Please refer to appendix for more ablations.
@@ 6 Conclusion
In this paper, we present CleanPose, the first solution that addresses the dataset biases in category-level pose estimation from the perspective of causal learning. Motivated by pivotal observation that humans can learn inherent causality beyond biases, we formulate the modeling of crucial causal variables and develop a causal inference framework in COPE task. Furthermore, we devise a residual knowledge distillation network to transfer unbiased semantics knowledge from 3D foundation model, providing comprehensive causal guidance to achieve unbiased estimation. Extensive experiments on challenging benchmarks demonstrate that CleanPose can significantly improve performance, showing the effectiveness of our method.
Acknowledgments
This paper is supported by the National Natural Science Foundation of China under Grants (62233013, 62173248, 62333017, 624B2105).
@@ Acknowledgments
A More Loss Function Details
The backbone of our method is based on AG-Pose [22]. In addition to $\mathcal{L}_{pose}$ (Eq. (21)), there are some additional loss functions to balance keypoints selection and pose prediction. First, to encourage the keypoints to focus on different parts, the diversity loss $\mathcal{L}_{div}$ is used to force the detected keypoints to be away from each other, in detail:
@@ A More Loss Function Details
Hence, the complete form of overall loss (Eq. (22)) is as follows:
@@
B More Details of Using ULIP-2
ULIP-2 [47] is a large-scale 3D foundation model with strong perception capabilities for the point cloud modality. It offers multiple pre-trained versions of point cloud encoders. In our model, there are two key steps that involve the use of different pre-trained encoders of ULIP-2. (i) In the knowledge distillation process, we leverage the pre-trained PointBert [50], which achieves the best zero-shot classification performance across all versions. Therefore, it can provide comprehensive category knowledge guidance for our model. In the ablation study Tab. 5d, we also compared it with PointNet++ [35], which is more similar in architecture to our model. The experimental results demonstrate that our distillation method focuses more on category knowledge rather than feature similarity. (ii) However, during the initial construction of the dynamic queue, we use the pre-trained PointNet++ [35], as the front-door adjustment primarily focuses on the differences between samples. We aim to avoid introducing confounders due to feature discrepancies from different encoders. The additional ablation study results in Tab. 2(b) also support our analysis.
C Additional Ablations
@@
Table S1: Additional ablation studies on some hyper-parameters. Default settings are colored in gray.
Effect of different sampling quantities $N_{s}$. In Sec. 4.2 of main manuscript, we sample $N_{s}$ features for the specific network design to perform front-door adjustment. We study the influence with response to different sampling quantities $N_{s}$ in Tab. 1(b). It can be found that a large $N_{s}$ leads to a slight performance degradation. We speculate that a larger sample size may introduce noise and redundant information that affects key features in causal inference. An appropriate sample size can balance valid and redundant information, prompting the model to focus on learning more representative causal correlations. The results demonstrate that $N_{s}=12$ yields the most significant performance gains.
Effect of different 3D encoder for initial construction of the queue. Tab. 2(b) ablates the different 3D encoders of ULIP-2 [47] for initial construction of the dynamic queue. The results exhibit that using PointNet++ [35] yields the most performance gains. As mentioned in Sec. B, the dynamic queue is utilized in the cross-attention phase of front-door adjustment, thus primarily focusing on the differences between samples. Using encoders with similar architectures helps avoid introducing extra confounders.
@@
PointNet++[35] 61.7 67.6 78.3 86.3
PointMLP[30] 58.1 65.8 76.3 85.2
PointBert[50] 58.8 65.6 77.9 86.4
@@
Random 61.7 67.6 78.3 86.3
K-means 58.5 65.5 78.0 86.5
K-means (simi) 58.7 66.2 77.7 86.0
@@
Table S2: Additional ablation studies on confounders queue. Default settings are colored in gray.
Figure S1: Illustration of different feature selectors in ablations Tab. 2(c).
Various feature selection strategies. In Sec. 4.2 of main manuscript, we randomly sample $N_{s}$ features from queue by default, as shown in Fig. S1(a). Optionally, we can first use K-means to cluster the features of the queue, and then select features from each cluster to form $\mathcal{F}_{samp}$, as illustrated in Fig. S1(b). For fair comparison, the number of clusters is set equal to $N_{s}$. We investigate the impact of these two feature selector in Tab. 2(c). As shown in the table, we observe that K-means-based feature sampling strategy shows a decline in performance on strict metrics (5°2cm and 5°5cm). We argue that k-means, which clusters by Euclidean distance, may lose important boundary information, thus affecting the model performance. Moreover, k-means clustering needs to be performed after each queue update, which increases computational load and training costs. Therefore, we added a comparative experiment using similarity-based updates, denoted as ‘K-means (simi)’, where (simi) refers to the ‘Similarity’ defined in Tab. 5a of main text.
Category $IoU_{25}^{*}$$\uparrow$ $IoU_{50}^{*}$$\uparrow$ $IoU_{75}^{*}$$\uparrow$ 5°2$\uparrow$ 5°5$\uparrow$ 10°2$\uparrow$ 10°5$\uparrow$
bottle 51.3 49.4 37.1 75.7 81.7 79.9 87.8
bowl 100.0 100.0 93.8 93.3 98.2 95.0 99.9
camera 90.9 83.5 39.9 2.8 3.2 33.9 40.5
can 71.3 71.1 43.2 84.2 85.9 96.9 98.6
laptop 86.3 84.0 76.1 69.2 90.9 71.9 98.5
mug 99.6 99.4 86.1 45.2 45.6 91.9 91.9
Average 83.3 81.2 62.7 61.7 67.6 78.3 86.3
Table S3: Category-wise evaluation of CleanPose on REAL275 dataset. ‘*’ denotes CATRE [28] IoU metrics.
@@
Seed 1 42 500 1k 1w 10w $\sigma^{2}\downarrow$
5°2cm 61.4 61.5 61.7 61.4 61.3 61.7 0.03
5°5cm 67.2 67.3 67.5 67.2 67.1 67.6 0.04
Table S5: Effect of different sampling seed during inference.
In this case, clustering is only performed once during the initial training. However, experimental results also show that such strategy leads to further performance degradation as one clustering loses the diversity of features.
Effect of different sampling seed during inference. We have conducted additional experiments with 6 different random seeds, as shown in Tab. S5. The computed variances $\sigma^{2}$ for metrics demonstrate stable performance across different random seeds, indicating the robustness and reliability of our method.
D More Experimental Results
@@
We report category-wise results of REAL275 [42] in Tab. S3. Since there is a small mistake in the original evaluation code of NOCS [42] for the 3D IoU metrics, we present more reasonable CATRE [28] metrics following [57, 26, 3]. Further, more detailed results of HouseCat6D [15] are shown in Tab. S4. As for more restricted metric $IoU_{75}$, our method also demonstrates state-of-the-art performance (53.9%), further validating the effectiveness of CleanPose in 3D IoU evaluation.
[…]
1 AG-Pose ViT-S/14 223M 61.3 57.0 51.3 35
2 ours ViT-S/14 246M 62.7 61.7 51.8 33
3 ours ViT-S/14 246M 61.8 58.1 51.6 33
4 AG-Pose ResNet18 220M 60.9 56.2 51.2 35
5 ours ResNet18 243M 62.3 60.3 51.6 33
[…]
Moreover, in category-wise validation on $IoU_{25}$ and $IoU_{50}$, our approach obtains state-of-the-art (e.g., Can, Cup, Glass and Shoe) or competitive results across all categories. It is worth mentioning that our method exhibits more stable performance on these two metrics. For instance, compared to the current sota method AG-Pose [22] in the Box category, our method achieves the best performance (28.7%) on $IoU_{50}$ metric when both obtain competitive results on $IoU_{25}$ metric, with a significant reduction of the AG-Pose (13.9%).
Figure S2: Effect of debiasing. Illustrated by Yaw angle distributions of the mug category.
E Detailed Comparison with SOTA method
We follow the domain consensus to report the metric accuracy in main manuscript. Moreover, we add more terms, e.g. visual encoder type, inference latency (FPS), in Tab. S6 for comprehensive comparison. As confirmed in (#1) and (#2) of Tab. S6, the front-door adjustment only increases the number of parameters by 10% (246M vs. 223M), while the running time remains nearly unchanged (33 vs. 35 in FPS). To ensure a fair comparison, we also replace the front-door module with MLPs that have the same number of parameters (#3). The results further demonstrate the superior effectiveness of causal learning. What’s more, we include additional results with ResNet18. Specifically, our method still outperforms AG-Pose with ResNet18 setting (#4 vs. #5), further supporting the efficacy of our approach.
F Effect of Debiasing.
We evaluate the debiasing effect via rotation distributions. As shown in Fig. S2, the predictions of baseline model are clearly biased toward the training set distributions, while the debiased model primarily unaffected.
G Limitation and Broader Impact